OpenStudy (anonymous):

A ball is thrown from a cliff with an intitial velocity of 92ft/sec. The height h of the object t seconds after it is released is given by the quadratic equation: h= -16t^2+92t+24

OpenStudy (sidsiddhartha):

v = u + at = 92 - 32t now v=0 t=2.875 now h=1/2at^2=0.5*32*(2.875)^2=132.25

OpenStudy (anonymous):

A ball is thrown from a cliff with an intitial velocity of 92ft/sec. The height h of the object t seconds after it is released is given by the quadratic equation: h= -16t^2+92t+24